trying to improve these answers or correct any that are incorrect.

Fresh Air Anti-Pollution Company is suffering declining sales of its principal product, non-biodegradable plastic cartons. The president, Tyler Weber, instructs his controller, Robin Cain, to lengthen asset lives to reduce depreciation expense. A processing line of automated plastic extruding equipment, purchased for $3.5 million in January 2014, was originally estimated to have a useful life of 8 years and a salvage value of $400,000. Depreciation has been recorded for 2 years on that basis. Tyler wants the estimated life changed to 12 years total and the straight-line method continued. Robin is hesitant to make the change, believing it is unethical to increase net income in this manner. Tyler says, “Hey, the life is only an estimate, and I've heard that our competition uses a 12-year life on their production equipment.”

1. Who are the stakeholders in this situation?

stockholders, potential investors, company managers and external users like creditors, bankers, government and taxing authorities

2. Is the proposed change in asset life unethical, or is it simply a good business practice by an astute president?

The practice is unethical. Since sales of their principal product are declining, the life of the equipment will likely become less not more, because sales of the product it is producing is likely to continue to decline. I also think it is unethical for the president to give accounting directives to the controller. The controller should be handling financial reporting independently without input from the president according to GAAP.

3. What is the effect of Tyler's proposed change on income before taxes in the year of change?

The depreciation expense for this asset at the end of the current year will be $272,500 instead of $387,500. This increases net income by $115,000.
Answer #1

1. Since Stakeholder is any person,organisation,group or society that has a stake in the business therefore, in the present case the company stakeholders are stockholders i.e the owners of the company,potential investors,managers of the company,employees,creditors,debtors,bankers,government and taxing authorities.

2.Estimate of Life of an asset is an accounting estimate by management and it is not unethical to revise the estimate.However, Changes in Accounting Estimates must be accounted for prospectively in the financial statements, i.e. the effects of the change must be incorporated in the accounting period in which the estimates are revised.

3. The depreciation recorded for 2 years on the basis of 8 year life will not be changed.
Depreciation on the basis of 8 years : (3500000-400000)/8=387500 per year
Value of Asset at end of 2 years: 3500000-387500-387500=2725000
Depreciation on the basis 12 year: 2725000/10=272500
Net increase in Income by way of decrease in depreciation: 378500-272500=115000